# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 8.6 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (Mandlekar et al. (2023)) and DexMimicGen (Jiang et al. (2024c)) change the positions and orientations of objects, while CACTI (Mandi et al. (2022)) adds new, artificial objects to the scene. However, the reliability of data augmentation still needs validation. For example, MimicGen (Mandlekar et al. (2023)) filters data generation attempts based on task success. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@ 9.1 What is the framework for general manipulation?
(1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates, precision and safety for specific tasks (Pang et al. (2023)). Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework proposed in this paper is based on the development of learning-based methods and the definition of general manipulation. The framework of brain-like cognitive research should also be explored.
@@ 9.2 What kind of learning capability should a general manipulation framework possess?
As an intelligent robot for general manipulation, it is inevitable that one cannot learn all the skills of an open-world during offline development, hence possessing a certain learning capability is necessary. Within the framework of this paper, a module of corrective instruction is introduced, enabling the robot to rectify its actions. These corrective demonstrations are incorporated into the manipulation dataset and used to improve the policy offline through fine-tuning. However, this approach still focuses on learning old task skills and cannot acquire new ones.
@@
The model of Policy should possess the capability of interactive, few-shot, continue, online learning to acquire a new skill and reinforce the policy’s mastery of the newly learned skill through corrective instruction offline. Interactive refers to the ability to learn through human demonstration or by observing instructional videos. Learning through demonstration often requires physical control or teleoperation, which is less natural. Learning through observation of instructional videos aligns better with human learning patterns. However, when humans learn from teachers, they often do not predict the teacher’s trajectory but rather understand the high-level description of the actions, akin to VLaMP (Patel et al. (2023)). Few-shot continue learning enables the robot to learn new skills with minimal demonstrations without forgetting previously learned skills. Online learning entails processing observed data instantly and enabling the model to learn as quickly as possible.
@@ 9.4 How to use internet-scale video data for RFMs?
As for what information from video dataset can be used, there are six main types of information to convert from video datasets: (1) Pose, such as capturing human hand poses and retargeting them to dexterous hand poses (Shaw et al. (2023); Qin et al. (2022)). (2) Affordance, including grasp locations on objects and post-grasp waypoints (Mendonca et al. (2023)). (3) Motion information, explicitly includes keypoints trajectories of objects and human hand during actions (Xiong et al. (2021); Yuan et al. (2024); Wen et al. (2023b)) and implicitly includes using VQ-VAE (Van Den Oord et al. (2017)) to generate a codebook for latent delta action (Ye et al. (2024)). (4) Environment transition dynamic information, such as predicting hindsight images after completing the current action (Wu et al. (2023a); Cheang et al. (2024); Yang et al. (2023b)). (5) Semantic information, such as descriptions of current task steps (Wang et al. (2024a)) and task instruction (Jain et al. (2024)). (6) Spatial and texture information, such as MVP (Radosavovic et al. (2023)) suggests using masked autoencoding (He et al. (2022)) for improving visual reconstruction.
@@
As for how to extract these useful information, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (Shaw et al. (2023)), affordance (Mendonca et al. (2023)), key points trajectory (Wen et al. (2023b)), latent action (Ye et al. (2024)), mask and bounding boxes (Shan et al. (2020)). When adding various labels to the video dataset, different training objectives can be used to extract features from the video dataset, such as MAE (Radosavovic et al. (2023)), contrastive learning (Ma et al. (2022)), time contrastive learning (Ma et al. (2023a)), temporal-difference learning (Bhateja et al. (2023)), video prediction objective(Du et al. (2024)), affordance prediction objective(Mendonca et al. (2023)), video-language alignment objective (Nair et al. (2022)), action motion objective(Yuan et al. (2024)) or combinations of these objectives (Karamcheti et al. (2023); Zhou et al. (2021)).
@@
As for how to utilize the extracted information to enhance or train robotic foundation models, the current robotic foundation models primarily use two learning methods: imitation learning and reinforcement learning. Therefore, the discussion on the third issue focuses on leveraging prior knowledge from video datasets in these two methods. As for imitation learning, when the robotic foundation model outputs pose and the video dataset annoated with pose label, the video dataset can be directly used as training data for the robotic foundation model (Shaw et al. (2023); Qin et al. (2022); Kareer et al. (2024)). When leveraging affordance information, motion information, environment transition dynamics information, semantic information, spatial and texture information, it is essential to employ GMM & CEM (Mendonca et al. (2023)), Inverse Dynamic Model (IDM) (Du et al. (2024); Ye et al. (2024); Wen et al. (2023b)), and Decoder (Wang et al. (2023a); Xiao et al. (2022b); Cheang et al. (2024); Wu et al. (2023a)) to transform these information into actions. Compared to other types of information, using semantic information treats the video as task instruction rather than observation (Jain et al. (2024); Shah et al. (2023); Jang et al. (2022)). At the same time, semantic information can also be used to organize tasks into a hierarchy of skills (Wang et al. (2024a)).
@@
As for reinforcement learning, the environment transition dynamics can be used as a transition model (Yang et al. (2023b)). The encoder, trained on a video dataset with various objectives, can measure the distance between cross-embodiment actions, which then serves as the reward function or value function (Bhateja et al. (2023)). For example, Guzey et al. (2024) and Xiong et al. (2021) use key points motion information to construct the encoder, which serves as the reward function for reinforcement learning. Since distribution shifts exist between cross-embodiment actions, AVID (Smith et al. (2019)) and LbW (Xiong et al. (2021)) translate human action images from videos into robot action images. However, this translation is limited to the pixel level.
@@
Current research focuses on different types of information in video datasets. The methods for extracting and using this information vary. It is important to consider which information from video datasets should be robustly applied to robotic foundation models. Video is similar to how humans perceive the world. Humans can improve their skills by watching experts. Similarly, using video datasets to construct a reinforcement learning from human feedback (RLHF) system in robotic foundation models is worth exploring (Luo et al. (2024)).
@@ 9.5 How to uses foundation models for post-conditions detection and post-hoc correction?
The current data collection mostly focuses on gathering successful task execution data, ignoring the collection of data related to failed task executions. However, if data on failed task executions are collected and annotated with corresponding error reasons, it would be possible to train a model to both determine task execution success and analyze the reasons for task execution failure. AHA (Duan et al. (2024)) trains a VLM to assess failures and output the reasons for these failures. However, the categories of failure modes are still limited, and it cannot output more open-ended failures, such as collaboration errors in dual-arm tasks. Many current studies use internet video data to improve the generalization of policies. Exploring the use of internet video data to enhance post-condition detection and employing multimodal perception to more accurately identify the reasons for failures is a promising direction. Post-hoc correction could then generate corrective action sequences based on the reasons for task execution failure and the task objectives, which would be handed over to a policy to generate corresponding corrective actions.
@@ 9.6 How to use foundation models for End-effector Design?
Currently, there are two primary approaches to designing end-effector. The first approach customizes the end-effector for specific tasks. The second approach makes the multi-fingered end-effector resemble a human hand. The end-effector designed with the first approach is usually easier to control because it has fewer degrees of freedom compared to the end-effector designed with the second approach. In Billard and Kragic (2019), dexterity is divided into two types: extrinsic dexterity and intrinsic dexterity. Extrinsic dexterity involves using external support, such as friction, gravity, and contact surfaces, to compensate for the lack of degrees of freedom. Intrinsic dexterity refers to the hand’s ability to manipulate objects using its own degrees of freedom. Therefore, the first approach still has certain limitations for general manipulation.
The first approach requires manual design, extensive testing, and continual adjustments. In Stella et al. (2023), LLMs are used for designing end-effector. However, this area is still in its early exploration stages. Using LLMs for end-effector design generates text descriptions, which still need to be manually translated into designs. This process is not fully automated. If we could develop modules for rotational and translational joints, and use something like protein structure prediction networks (Jumper et al. (2021)), training a foundation model to output graph including these joints could help reduce the challenges of manual design. As for the second approach, the human hand has many sensors and actuators. This makes it nearly impossible to design a robotic hand that closely resembles the human hand. Therefore, it’s essential to design the sensors and actuators carefully.
@@
RLBench (James et al. (2020)) Rigid/Articulated/Deformable 28 Table-top 100 90 CoppeliaSim
Behavior-1K(Li et al. (2023a)) Rigid/Articulated/Deformable 9318 50 1000 - Omniverse
VirtualHome (Puig et al. (2018)) Rigid/Articulated/Deformable 1138 50 8014 5193 Unity
RoboCasa (Nasiriany et al. (2024)) Rigid/Articulated 2509 10 100 $\times$ 100K+ Mujoco
LIBERO-1K(Liu et al. (2024a)) Rigid/Articulated 67 Table-top 130 6500 Mujoco
Robosuite (Zhu et al. (2020)) Rigid/Articulated 20 Table-top 9 $\times$ - Mujoco
Sapien(Xiang et al. (2020)) Rigid/Articulated 2346 Table-top 5 $\times$ - Sapien(NVIDIA PhysX+OpenGL)
Maniskill2 (Gu et al. (2023)) Rigid/Articulated/Deformable 2144 - 20 $\times$ 30K+ Sapien
CALVIN(Mees et al. (2022)) Rigid/Articulated 28 4 34 40M Pybullet
@@ 9.7 How to use foundation models for dexterous manipulation?
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Although model-based hard code method (Zhu et al. (2024)) can collect data on dexterous manipulation, they still require data analysis such as mutual information (Hejna et al. (2025)) and entropy (Zhu et al. (2024)) to assess the quality of the data. Additionally, for multi-scenario and multi-task data collection, teleoperation methods are less dependent on algorithm performance compared to model-based hard code methods. However online teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the real-robot (Wang et al. (2024c)).
Two main learning-based methods for dexterous manipulation are imitation learning (Ze et al. (2024a)) and reinforcement learning (Ma et al. (2023b)). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (Chi et al. (2023)) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. It can also use an existing RFMs for fine-tuning (in Sec. 7). Fine-tuning with a RFMs allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (Brohan et al. (2023)).
Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (Ye et al. (2023a)) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (Gupta et al. (2021)) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
Current learning methods each have their strengths and weaknesses (Zhang et al. (2024a)). Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the Appendix.D.
@@ 9.8 How to use foundation models for whole-body control?
The above discussion primarily focuses on the contact between the end-effector and the object. However, whole-body control is still needed in dexterous manipulation. For example, in a polishing robot, force-position hybrid control of the robotic arm is often required to manage the trajectory of contact points and forces/torques. Mobile manipulation is essential for dexterous manipulation reachability. This idea is inspired by how humans handle objects. For example, when playing badminton, people use their waists, shoulders, elbows, and wrists together to hit the shuttlecock further. This aspect is often overlooked by current foundational models for manipulation. Although LEO (Huang et al. (2023a)) can provide poses for both navigation and manipulation, it still does not address the synchronization issue between the two.
For whole-body control, the focus is on low-level control issues. A straightforward idea is to expand the action space of the policy model to include all joints of the robot. However, as the output dimensions increase, end-to-end training methods are more likely to diverge. Therefore, most current models output cartesian space poses and force/torques. These outputs are then optimized and converted into position or torque for each joint through a post-processing module (Haviland and Corke (2021)). To address end-to-end whole-body control issues, principal research is needed to facilitate network training and deployment.

### v7

@@ 8.6 Dataset Augmentation
Current data augmentation can be mainly divided into scene-level and object-level. Scene-level refers to changing the layout of objects in the scene. For example, MimicGen (mandlekar2023mimicgen) and DexMimicGen (jiang2024dexmimicgen) change the positions and orientations of objects, while CACTI (mandi2022cacti) adds new, artificial objects to the scene. However, the reliability of data augmentation still needs validation. For example, MimicGen (mandlekar2023mimicgen) filters data generation attempts based on task success. Current foundation models for dataset augmentation methods primarily operate at the object level. The main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s texture. GenAug (chen2023genaug) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (yu2023scaling) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
@@ 9.1 What is the framework for general manipulation?
(1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates, precision and safety for specific tasks (pang2023global). Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework proposed in this paper is based on the development of learning-based methods and the definition of general manipulation. The framework of brain-like cognitive research should also be explored.
@@ 9.2 What kind of learning capability should a general manipulation framework possess?
As an intelligent robot for general manipulation, it is inevitable that one cannot learn all the skills of an open-world during offline development, hence possessing a certain learning capability is necessary (wang2024imperative). Within the framework of this paper, a module of corrective instruction is introduced, enabling the robot to rectify its actions. These corrective demonstrations are incorporated into the manipulation dataset and used to improve the policy offline through fine-tuning. However, this approach still focuses on learning old task skills and cannot acquire new ones.
@@
The model of Policy should possess the capability of interactive, few-shot, continue, online learning to acquire a new skill and reinforce the policy’s mastery of the newly learned skill through corrective instruction offline. Interactive refers to the ability to learn through human demonstration or by observing instructional videos. Learning through demonstration often requires physical control or teleoperation, which is less natural. Learning through observation of instructional videos aligns better with human learning patterns. However, when humans learn from teachers, they often do not predict the teacher’s trajectory but rather understand the high-level description of the actions, akin to VLaMP (patel2023pretrained). Few-shot continue learning enables the robot to learn new skills with minimal demonstrations without forgetting previously learned skills. Online learning entails processing observed data instantly and enabling the model to learn as quickly as possible.
@@ 9.4 How to use internet-scale video data for RFMs?
As for what information from video dataset can be used, there are six main types of information to convert from video datasets: (1) Pose, such as capturing human hand poses and retargeting them to dexterous hand poses (shaw2023videodex; qin2022dexmv). (2) Affordance, including grasp locations on objects and post-grasp waypoints (mendonca2023structured). (3) Motion information, explicitly includes keypoints trajectories of objects and human hand during actions (xiong2021learning; yuan2024general; wen2023any) and implicitly includes using VQ-VAE (van2017neural) to generate a codebook for latent delta action (ye2024latent). (4) Environment transition dynamic information, such as predicting hindsight images after completing the current action (wu2023unleashing; cheang2024gr; yang2023learning). (5) Semantic information, such as descriptions of current task steps (wang2024vlm) and task instruction (jain2024vid2robot). (6) Spatial and texture information, such as MVP (radosavovic2023real) suggests using masked autoencoding (he2022masked) for improving visual reconstruction.
@@
As for how to extract these useful information, various off-the-shelf models can be used to annotate the current video with additional labels, such as pose (shaw2023videodex), affordance (mendonca2023structured), key points trajectory (wen2023any), latent action (ye2024latent), mask and bounding boxes (shan2020understanding). When adding various labels to the video dataset, different training objectives can be used to extract features from the video dataset, such as MAE (radosavovic2023real), contrastive learning (ma2022vip), time contrastive learning (ma2023liv), temporal-difference learning (bhateja2023robotic), video prediction objective(du2024learning), affordance prediction objective(mendonca2023structured), video-language alignment objective (nair2022r3m), action motion objective(yuan2024general) or combinations of these objectives (karamcheti2023language; zhou2021ibot).
@@
As for how to utilize the extracted information to enhance or train robotic foundation models, the current robotic foundation models primarily use two learning methods: imitation learning and reinforcement learning. Therefore, the discussion on the third issue focuses on leveraging prior knowledge from video datasets in these two methods. As for imitation learning, when the robotic foundation model outputs pose and the video dataset annoated with pose label, the video dataset can be directly used as training data for the robotic foundation model (shaw2023videodex; qin2022dexmv; kareer2024egomimic). When leveraging affordance information, motion information, environment transition dynamics information, semantic information, spatial and texture information, it is essential to employ GMM & CEM (mendonca2023structured), Inverse Dynamic Model (IDM) (du2024learning; ye2024latent; wen2023any), and Decoder (wang2023mimicplay; xiao2022masked; cheang2024gr; wu2023unleashing) to transform these information into actions. Compared to other types of information, using semantic information treats the video as task instruction rather than observation (jain2024vid2robot; shah2023mutex; jang2022bc). At the same time, semantic information can also be used to organize tasks into a hierarchy of skills (wang2024vlm).
@@
As for reinforcement learning, the environment transition dynamics can be used as a transition model (yang2023learning). The encoder, trained on a video dataset with various objectives, can measure the distance between cross-embodiment actions, which then serves as the reward function or value function (bhateja2023robotic). For example, guzey2024bridging and xiong2021learning use key points motion information to construct the encoder, which serves as the reward function for reinforcement learning. Since distribution shifts exist between cross-embodiment actions, AVID (smith2019avid) and LbW (xiong2021learning) translate human action images from videos into robot action images. However, this translation is limited to the pixel level.
@@
Current research focuses on different types of information in video datasets. The methods for extracting and using this information vary. It is important to consider which information from video datasets should be robustly applied to robotic foundation models. Video is similar to how humans perceive the world. Humans can improve their skills by watching experts. Similarly, using video datasets to construct a reinforcement learning from human feedback (RLHF) system in robotic foundation models is worth exploring (luo2024serl).
@@ 9.5 How to uses foundation models for post-conditions detection and post-hoc correction?
The current data collection mostly focuses on gathering successful task execution data, ignoring the collection of data related to failed task executions. However, if data on failed task executions are collected and annotated with corresponding error reasons, it would be possible to train a model to both determine task execution success and analyze the reasons for task execution failure. AHA (duan2024aha) trains a VLM to assess failures and output the reasons for these failures. However, the categories of failure modes are still limited, and it cannot output more open-ended failures, such as collaboration errors in dual-arm tasks. Many current studies use internet video data to improve the generalization of policies. Exploring the use of internet video data to enhance post-condition detection and employing multimodal perception to more accurately identify the reasons for failures is a promising direction. Post-hoc correction could then generate corrective action sequences based on the reasons for task execution failure and the task objectives, which would be handed over to a policy to generate corresponding corrective actions.
@@ 9.6 How to use foundation models for End-effector Design?
Currently, there are two primary approaches to designing end-effector. The first approach customizes the end-effector for specific tasks. The second approach makes the multi-fingered end-effector resemble a human hand. The end-effector designed with the first approach is usually easier to control because it has fewer degrees of freedom compared to the end-effector designed with the second approach. In billard2019trends, dexterity is divided into two types: extrinsic dexterity and intrinsic dexterity. Extrinsic dexterity involves using external support, such as friction, gravity, and contact surfaces, to compensate for the lack of degrees of freedom. Intrinsic dexterity refers to the hand’s ability to manipulate objects using its own degrees of freedom. Therefore, the first approach still has certain limitations for general manipulation.
The first approach requires manual design, extensive testing, and continual adjustments. In stella2023can, LLMs are used for designing end-effector. However, this area is still in its early exploration stages. Using LLMs for end-effector design generates text descriptions, which still need to be manually translated into designs. This process is not fully automated. If we could develop modules for rotational and translational joints, and use something like protein structure prediction networks (jumper2021highly), training a foundation model to output graph including these joints could help reduce the challenges of manual design. As for the second approach, the human hand has many sensors and actuators. This makes it nearly impossible to design a robotic hand that closely resembles the human hand. Therefore, it’s essential to design the sensors and actuators carefully.
@@
RLBench (james2020rlbench) Rigid/Articulated/Deformable 28 Table-top 100 90 CoppeliaSim
Behavior-1K(li2023behavior) Rigid/Articulated/Deformable 9318 50 1000 - Omniverse
VirtualHome (puig2018virtualhome) Rigid/Articulated/Deformable 1138 50 8014 5193 Unity
RoboCasa (nasiriany2024robocasa) Rigid/Articulated 2509 10 100 $\times$ 100K+ Mujoco
LIBERO-1K(liu2024libero) Rigid/Articulated 67 Table-top 130 6500 Mujoco
Robosuite (zhu2020robosuite) Rigid/Articulated 20 Table-top 9 $\times$ - Mujoco
Sapien(xiang2020sapien) Rigid/Articulated 2346 Table-top 5 $\times$ - Sapien(NVIDIA PhysX+OpenGL)
Maniskill2 (gu2023maniskill2) Rigid/Articulated/Deformable 2144 - 20 $\times$ 30K+ Sapien
CALVIN(mees2022calvin) Rigid/Articulated 28 4 34 40M Pybullet
@@ 9.7 How to use foundation models for dexterous manipulation?
One major challenge in data collection for dexterous manipulation lies in gathering data from multi-fingered end-effectors. Although model-based hard code method (zhu2024should) can collect data on dexterous manipulation, they still require data analysis such as mutual information (hejna2025robot) and entropy (zhu2024should) to assess the quality of the data. Additionally, for multi-scenario and multi-task data collection, teleoperation methods are less dependent on algorithm performance compared to model-based hard code methods. However online teleoperation requires a real-robot system, which is not portable and cannot achieve in-the-wild data collection. Therefore, current mainstream research focuses on directly tracking human hand motions during manipulation without controlling the real-robot (wang2024dexcap).
Two main learning-based methods for dexterous manipulation are imitation learning (ze2024h) and reinforcement learning (ma2023eureka). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (chi2023diffusion) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. It can also use an existing RFMs for fine-tuning (in Sec. 7). Fine-tuning with a RFMs allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (rt2).
Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (ye2023foundation) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (gupta2021reset) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
Current learning methods each have their strengths and weaknesses (zhang2024hirt). Therefore, learning approaches for dexterous manipulation should integrate different methods. For example, diffusion policy can assist reinforcement learning in addressing high-dimensional action spaces issue, while reinforcement learning can help diffusion policy overcome issues with suboptimal and negative data. Additionally, the learning models should consider both inputs and outputs. The factors necessary for achieving dexterous manipulation are summarized in the Appendix.D.
@@ 9.8 How to use foundation models for whole-body control?
The above discussion primarily focuses on the contact between the end-effector and the object. However, whole-body control is still needed in dexterous manipulation. For example, in a polishing robot, force-position hybrid control of the robotic arm is often required to manage the trajectory of contact points and forces/torques. Mobile manipulation is essential for dexterous manipulation reachability. This idea is inspired by how humans handle objects. For example, when playing badminton, people use their waists, shoulders, elbows, and wrists together to hit the shuttlecock further. This aspect is often overlooked by current foundational models for manipulation. Although LEO (huang2023embodied) can provide poses for both navigation and manipulation, it still does not address the synchronization issue between the two.
For whole-body control, the focus is on low-level control issues. A straightforward idea is to expand the action space of the policy model to include all joints of the robot. However, as the output dimensions increase, end-to-end training methods are more likely to diverge. Therefore, most current models output cartesian space poses and force/torques. These outputs are then optimized and converted into position or torque for each joint through a post-processing module (haviland2021neo). To address end-to-end whole-body control issues, principal research is needed to facilitate network training and deployment.